a man loses rs. 20 by selling some toys at the rate of rs.3 per piece and gains rs. 30 if he sells them at rs. 3.25 per piece. find the number of pieces sold by him.
Answer by ankor@dixie-net.com(22740)


Let x = no. of pieces sold
Let c = his cost for these items
:
Write an equation for each statement
:
" a man loses rs. 20 by selling some toys at the rate of rs.3 per piece"
3x - c = -20
:
"and gains rs. 30 if he sells them at rs. 3.25 per piece."
3.25x - c = +30
:
Use elimination on the two equation
3.25x - c = 30
3x - c = -20
-------------------Subtraction eliminates c, find x
.25x = 50
x = 50/.25
x = 200 toys sold